# Active viscoelastic condensates provide controllable mechanical anchor points

### v1

@@ Introduction
Viscoelastic model of condensate growth
@@
The cell provides a finite reservoir of condensate material
@@
Internal reactions fuel condensate growth
@@
Here, $k_{+}$ is the rate of conversion of precursor material into scaffold material in the bulk of the condensate; $k_{-}$ is the rate of turnover of scaffold back into precursor; and ${q_{\mathrm{c}}(\bm{r})=q\delta(|\bm{r}-\bm{r}_{0}|-r_{\mathrm{c}})}$ quantifies the conversion of precursor into scaffold at the surface of the active core, which we parameterise as a rigid sphere with fixed radius $r_{\mathrm{c}}$ and position $\bm{r}_{0}$. In general, $k_{+}$, $k_{-}$, and $q$ need not be constant, but may depend on the strain in the scaffold network.
Material growth strains the scaffold
@@
where $\stackrel{{\scriptstyle\triangledown}}{{\mathbf{B}}}=\partial_{t}\mathbf{B}+\left[\bm{v}\cdot\nabla\right]\mathbf{B}-\left[\nabla\bm{v}\right]^{\intercal}\cdot\mathbf{B}-\mathbf{B}\cdot\left[\nabla\bm{v}\right]$ is the upper-convected derivative, $\mathbf{I}$ is the identity tensor, and $\tau$ is the characteristic relaxation time. The left side of equation (3) represents the rate of change of $\mathbf{B}$ with respect to a frame of reference both co-moving and co-rotating with the scaffold, as driven by the scaffold velocity $\bm{v}$, and the right side characterises the relaxation of $\mathbf{B}$ to the undeformed state ($\mathbf{B}=\mathbf{I}$) over timescale $\tau$. At the surface of the active core, $\mathbf{B}$ satisfies a homogeneous Neumann boundary condition. To calculate the elastic stress $\bm{\sigma}$ from the strain $\mathbf{B}$, we use a Neo-Hookean model [20],
@@
where $\Lambda$ is the mobility of scaffold material. In general, the osmotic stress is a function of $\phi_{S}$ that depends on the specific energetic interactions between the scaffold and its pore fluid. For a Flory–Huggins mixing energy, for example, the osmotic stress is isotropic, with diagonal components given by ${\Pi=-k_{\mathrm{B}}T\left[\phi_{S}+\log(1-\phi_{S})-\chi\phi_{S}^{2}\right]}$. Here, $k_{\mathrm{B}}T$ is the characteristic energy scale, and $\chi$ is the Flory interaction parameter. Variations in either $\phi_{S}$ or $\mathbf{B}$ (and hence $\bm{\sigma}$) can thus cause a change in the scaffold flux. In turn, the scaffold flux drives changes in both $\phi_{S}$ and $\mathbf{B}$ (via equations (1) and (3), respectively), therefore providing a two-way coupling between the reaction–diffusion kinetics discussed above and viscoelastic deformation of the condensate.
Elastic stresses favour a uniform scaffold
To illustrate the key features of our system, we first conduct detailed simulations of the above equations by using a phase-field formulation, in which the boundary between the condensate interior and exterior is approximated as a diffuse interface. In this framework, we treat $\phi_{S}$ as a phase-field order parameter that smoothly interpolates between the value of $\phi_{S}$ inside the condensate and zero outside the condensate. We describe the transport of precursor material in the cell by the reaction–diffusion equation ${\partial_{t}\phi_{P}=\nabla\left[D\cdot\nabla\phi_{P}\right]-s}$, where $D$ is a diffusion coefficient that may depend on composition. Along with this evolution equation for $\phi_{P}$, equations (1-5) give a closed system of equations that can be solved to find $\phi_{S}(\bm{r},t)$ and $\mathbf{B}(\bm{r},t)$. Assuming spherical symmetry, we solve these equations via the finite element scheme detailed in the Supplementary Information.
Figure 2: Example phase-field simulation of condensate growth, showing how the spatial distribution of scaffold material (A) and elastic pressure (B) evolve over time (light to dark curves). The inset shows corresponding results for a viscous condensate ($K=0$). Results are plotted at evenly spaced times up to $k_{-}t=0.4$. The dashed black line shows the initial condition. Reaction parameters values are $k_{+}/k_{-}=400$ and $q/k_{-}r_{\textrm{c}}=100$, and the rescaled viscoelastic relaxation time is $k_{-}\tau=0.1$. Other parameter values related to the phase-field formulation are provided in the Supplementary Information.
@@
Integrated model describes growth dynamics
Informed by our phase-field simulations, we now derive a simplified model that explicitly describes the evolution of the condensate radius $R$. Integrating equation (1) over the volume of the condensate, we find
@@
Finally, we assume that precursor material diffuses sufficiently rapidly that we can approximate $\phi_{P}$ by its mean value $\bar{\phi}_{P}$. Similarly, we take the volume fraction of precursor material outside the condensate to be uniform and given by $\phi_{P}^{\textrm{out}}=\eta\bar{\phi}_{P}$, where $\eta$ quantifies any differential partitioning of precursor between the condensate interior and exterior. As a result, $\bar{\phi}_{P}$ can be calculated directly from $\bar{\phi}_{S}$ and the condensate volume $V=\frac{4}{3}\pi(R^{3}-r_{\mathrm{c}}^{3})$ via conservation of total condensate material. Equations (6-8), in addition to the radial and azimuthal normal components of equations (3) and (4), then provide a closed system of equations that describes the evolution of the condensate radius, $R(t)$, and the strain profile within the scaffold, $B_{rr}(r,t)$ and $B_{\theta\theta}(r,t)$. These equations comprise a coupled ODE–PDE system that can be solved numerically given appropriate initial conditions.
Bulk incorporation accelerates growth
@@
Scaffold growth induces large strains near the core
We now study the mechanics of condensate growth, highlighting how different modes of material incorporation induce qualitatively different states of deformation. In general, incorporation of new scaffold components drives an outward flux of material, quantified by $j_{r}$, that strains the existing scaffold network. In the absence of bulk incorporation ($k_{+}=0$), all new material is inserted at the surface of the active core and fluxes outwards from the condensate centre. Consequently, $j_{r}$ is largest near to the core, and decreases monotonically towards the outer boundary of the condensate (figure 4, left). This flux induces a radial compression ($B_{rr}<1$) and azimuthal tension ($B_{\theta\theta}>1$) in the scaffold, which reduce with distance $r$ from the centre of the condensate. In contrast, bulk reactions promote isotropic material incorporation, such that new scaffold components are inserted throughout the condensate volume, and $j_{r}$ increases almost linearly with $r$. The scaffold flux at the surface of the core $r=r_{\textrm{c}}$ is set by the rate of core incorporation. If $q=0$, then $j_{r}=0$ at $r=r_{\textrm{c}}$ (figure 4, centre). In this case, scaffold molecules initially situated at the core surface remain at the core surface during condensate growth, and so cannot be stretched in the azimuthal direction ($B_{\theta\theta}|_{r=r_{\textrm{c}}}=1$). In the radial direction, however, the outwards flux of condensate material stretches the scaffold and induces large tensile strains. The scaffold flux resulting from the combination of both core and bulk incorporation is as expected; $j_{r}$ increases linearly far away from the core, but is non-zero at the core surface (figure 4, right). Correspondingly, we find large strains near to the core (which may be either tensile or compressive depending on the magnitude of $q$), and uniform tensile strains away from the core.
@@
Internal elasticity restricts condensate growth
To investigate the impact of internal elasticity on condensate growth, we now couple scaffold density to elastic deformation, allowing $\bar{\phi}_{S}$ to vary with stress as described by equation (8). Figure 5A shows that the influence of elastic stress on growth depends strongly on the rate of stress relaxation, quantified by $\tau$, compared to the rate of material incorporation. In the limit $k_{+}\ll\tau^{-1}$, strains relax sufficiently fast that condensate growth is unaffected by internal stresses. However, for $k_{+}\gg\tau^{-1}$ strain relaxation is slow compared to the rate of material incorporation and growth is significantly inhibited. Here, we keep $k_{+}$ and $q$ constant, implying that new material is incorporated into the condensate at the same rate regardless of the stress state. Consequently, elastic stresses have only a small effect on the rate of precursor material consumption (figure 5B). This result highlights that the restricted growth is driven by increased deformation of the scaffold, rather than by reaction kinetics.
We explain the impact of elasticity on growth as follows: as new material is incorporated into the condensate, the scaffold volume fraction $\phi_{S}$, and hence the osmotic pressure, increase locally. To reduce osmotic pressure and obtain the preferred volume fraction $\phi_{S}^{0}$, the condensate expands, with solvent swelling the scaffold. This expansion stretches the scaffold, inducing a large tensile stress. This stress then acts against the osmotic pressure, resisting the increase in tension associated with further swelling. As a result, the condensate cannot expand fully, and remains at a smaller size, with $\bar{\phi}_{S}>\phi_{S}^{0}$. Over time, stress relaxation allows the osmotic pressure to drive further scaffold expansion, until eventually $\bar{\phi}_{S}=\phi_{S}^{0}$ and the condensate reaches its final (unstrained) volume. Faster stress relaxation (smaller $\tau$, corresponding to a more liquid-like rheology) therefore facilitates faster condensate growth.
@@
Viscoelastic material properties impart mechanical strength
@@
Centrosomes as growing viscoelastic materials
Figure 6: (A) Phase space showing the quality of fit between simulation results and experimental growth curves for different elastic modulus $K$ and relaxation time $\tau$. The white curve separates parameter values for which an acceptable fit is possible, and for which it is not. (B) Normalised change in centrosome radius $\Delta R$ induced by a mitotic stress for various $K$ and $\tau$. The white curve separates parameter values for which scaffold expansion is acceptable, and for which it is not. (C) Combining results from panels (B) and (C) provides constraints on the material properties that allow both rapid growth and mechanical strength. (D) Proportion of total scaffold material that is incorporated at the centriole as a function of base incorporation rate $q_{0}$. Light to dark lines show simulations for increasing $\beta$ from $\beta=10^{-5}$ to $\beta=0.1$, which characterises how strongly reactions depend on strain. Other parameter values are the same as in figure 5. We convert our results to dimensional units by setting $k_{-}=10^{-3}\,\textrm{s}$, $r_{\textrm{c}}=75\,\textrm{nm}$, and $\alpha=0.1\,\textrm{Pa}^{-1}$.
Centrosomes are membrane-less organelles that act as the main microtubule organising centre in metazoans. Centrosomes consist of a well-defined core (a pair of centrioles), which is surrounded by an amorphous, mesh-like structure known as the pericentriolar material (PCM). The PCM comprises a porous scaffold network permeated by both solvent and client molecules [21]. Recent experiments have shown that the PCM in C. elegans centrosomes exhibits a characteristic pore size [22], and displays a viscoelastic rheology [23]. During mitosis, centrosomes mature rapidly, with the PCM dramatically increasing in size. Incorporation of new scaffold material is driven by phosphorylation-dependent conversion of precursor molecules into an assembly-competent form within the pre-existing PCM [24, 25, 26]. As centrosomes grow, microtubules nucleate within the PCM, extending outwards into the surrounding cell. These microtubules utilise centrosomes as mechanical anchor points for transducing mechanical forces that organise the mitotic spindle [27]. Current models of centrosome structure, however, are unable to reconcile the twin requirements of rapid (fluid-like) PCM assembly and (solid-like) mechanical strength [9, 10]. To demonstrate the ability of our modelling framework to explain this dichotomy, we compare the predictions of our model to previous experimental data for centrosomes in C. elegans embryos [28, 29, 30].
@@
Liquid-like centrosomes permit rapid scaffold growth
@@
During centrosome maturation, the PCM scaffold undergoes rapid expansion as new scaffold proteins (e.g., SPD-5 in the case of C. elegans) are incorporated. However, scaffold incorporation also drives large strains within the PCM, which can restrict growth (figure 5). Comparing experimental growth curves for centrosomes from embryos at the 1, 2 and 4 cell stages [28] with model predictions, we can constrain the mechanical parameters that allow sufficiently rapid growth (figure 6A). To determine permissible material properties, we fit our model to experimental data by varying $k_{+}$ and $\bar{\phi}$ for given values of the relaxation time $\tau$ and elastic modulus $K$. We keep the rate of centriolar incorporation $q$ constant throughout the fit, since it has a very limited effect on the overall growth curves. Quantifying the goodness of fit for each combination of material parameters by the mean squared error (figure 6A), we then identify ‘acceptable’ and ‘unacceptable’ regions of the parameter space for $\tau$ and $K$ spanning several orders of magnitude. We note that because elastic effects enter our model in combination with the phenomenological parameter $\alpha$, the dimensional values of $K$ presented here can be rescaled without affecting our general conclusions. For large values of $\tau$ and $K$ (solid-like rheology), we find that model predictions are incompatible with experimental data. In this parameter regime, a good fit between experiment and theory is not possible, regardless of the fit parameters $k_{+}$ and $\bar{\phi}$, since elastic stresses limit the maximum rate of growth. On the other hand, for small values of $\tau$ and $K$ (liquid-like rheology), an excellent fit is possible.
Viscoelastic centrosomes can sustain mitotic forces
[…]
Stress-dependent reactions promote bulk incorporation

### v2

@@ I Introduction
II Viscoelastic model of condensate growth
@@
II.1 The cell provides a finite reservoir of condensate material
@@
II.2 Internal reactions fuel condensate growth
@@
Here, $k_{+}$ is the rate of conversion of precursor material into scaffold material in the bulk of the condensate; $k_{-}$ is the rate of turnover of scaffold back into precursor; and $q_{\mathrm{c}}(\bm{r})=q\delta(r-r_{\mathrm{c}})$} quantifies the conversion of precursor into scaffold at the surface of the active core, which we parameterise as a rigid sphere with fixed radius $r_{\mathrm{c}}$. In general, $_{+}, kk_{-}$, and $q$ need not be constant, but may depend on the strain in the scaffold network.
II.3 Material growth strains the scaffold
@@
where $\stackrel{{\scriptstyle\triangledown}}{{\mathbf{B}}}=\partial_{t}\mathbf{B}+\left[\bm{v}\cdot\nabla\right]\mathbf{B}-\left[\nabla\bm{v}\right]^{\intercal}\cdot\mathbf{B}-\mathbf{B}\cdot\left[\nabla\bm{v}\right]$ is the upper-convected derivative, $\mathbf{I}$ is the identity tensor, and $\tau$ is the characteristic relaxation time. The left side of (3) represents the rate of change of $\mathbf{B}$ with respect to a frame of reference both co-moving and co-rotating with the scaffold, as driven by the scaffold velocity $\bm{v}$, and the right side characterises the relaxation of $\mathbf{B}$ to the undeformed state ($\mathbf{B}=\mathbf{I}$) over timescale $\tau$. At the surface of the active core, $\mathbf{B}$ satisfies a homogeneous Neumann boundary condition. To calculate the elastic stress $\bm{\sigma}$ from the strain $\mathbf{B}$, we use a Neo-Hookean model [20],
@@
where $\Lambda$ is the mobility of scaffold material. In general, the osmotic stress is a function of $\phi_{S}$ that depends on the specific energetic interactions between the scaffold and its pore fluid. For a Flory–Huggins mixing energy, for example, the osmotic stress is isotropic, with diagonal components given by $\Pi=-k_{\mathrm{B}}T\left[\phi_{S}+\log(1-\phi_{S})-\chi\phi_{S}^{2}\right]$}. Here, $k_{\mathrm{B}}T$ is the characteristic energy scale, and $\chi$ is the Flory interaction parameter. Variations in either $\phi_{S}$ or $\mathbf{B}$ (and hence $\bm{\sigma}$) can thus cause a change in the scaffold flux. In turn, the scaffold flux drives changes in both $\phi_{S}$ and $\mathbf{B}$ (via (1) and (3), respectively), therefore providing a two-way coupling between the reaction–diffusion kinetics discussed above and viscoelastic deformation of the condensate.
II.4 Elastic stresses favour a uniform scaffold
To illustrate the key features of our system, we first conduct detailed simulations of the above equations by using a phase-field formulation, in which the boundary between the condensate interior and exterior is approximated as a diffuse interface. In this framework, we treat $\phi_{S}$ as a phase-field order parameter that smoothly interpolates between the value of $\phi_{S}$ inside the condensate and zero outside the condensate. We describe the transport of precursor material in the cell by the reaction–diffusion equation $\partial_{t}\phi_{P}=\nabla\left[D\cdot\nabla\phi_{P}\right]-s$}, where $D$ is a diffusion coefficient that may depend on composition. Along with this evolution equation for $\phi_{P}$, Eqs. (1-5) give a closed system of equations that can be solved to find $\phi_{S}(\bm{r},t)$ and $\mathbf{B}(\bm{r},t)$. Assuming spherical symmetry, we solve these via the finite element scheme detailed in the Supplementary Information.
Figure 2: Example phase-field simulation of condensate growth, showing how the spatial distribution of scaffold material (A) and elastic pressure (B) evolve over time (light to dark curves). The inset shows corresponding results for a viscous condensate ($K=0$). Results are plotted at evenly spaced times up to $k_{-}t=0.4$. The dashed black line shows the initial condition. Reaction parameter values are $k_{+}/k_{-}=400$ and $q/k_{-}r_{\textrm{c}}=100$, and the rescaled viscoelastic relaxation time is $k_{-}\tau=0.1$. Other parameter values related to the phase-field formulation are provided in the Supplementary Information.
@@ II.4 Elastic stresses favour a uniform scaffold
II.5 Integrated model describes growth dynamics
Informed by our phase-field simulations, we now derive a simplified model that explicitly describes the evolution of the condensate radius $R$. Integrating (1) over the volume of the condensate, we find
@@ II.5 Integrated model describes growth dynamics
Finally, we assume that precursor material diffuses sufficiently rapidly that we can approximate $\phi_{P}$ by its mean value $\bar{\phi}_{P}$. Similarly, we take the volume fraction of precursor material outside the condensate to be uniform and given by $\phi_{P}^{\textrm{out}}=\eta\bar{\phi}_{P}$, where $\eta$ quantifies any differential partitioning of precursor between the condensate interior and exterior. As a result, $\bar{\phi}_{P}$ can be calculated directly from $\bar{\phi}_{S}$ and the condensate volume $V=\frac{4}{3}\pi(R^{3}-r_{\mathrm{c}}^{3})$ via conservation of total condensate material. Eqs. (6-8), in addition to the radial and azimuthal normal components of (3) and (4), then provide a closed system of equations that describes the evolution of the condensate radius, $R(t)$, and the strain profile within the scaffold, $B_{rr}(r,t)$ and $B_{\theta\theta}(r,t)$. These equations comprise a coupled ODE–PDE system that can be solved numerically given appropriate initial conditions.
II.6 Bulk incorporation accelerates growth
@@
II.7 Scaffold growth induces large strains near the core
We now study the mechanics of condensate growth, highlighting how different modes of material incorporation induce qualitatively different states of deformation. In general, incorporation of new scaffold components drives an outward flux of material, quantified by $j_{r}$, that strains the existing scaffold network. In the absence of bulk incorporation ($k_{+}=0$), all new material is inserted at the surface of the active core and fluxes outwards from the condensate centre. Consequently, $j_{r}$ is largest near to the core, and decreases monotonically towards the outer boundary of the condensate (figure 4; left). This flux induces a radial compression ($B_{rr}<1$) and azimuthal tension ($B_{\theta\theta}>1$) in the scaffold, which reduce with distance $r$ from the centre of the condensate. In contrast, bulk reactions promote isotropic material incorporation, such that new scaffold components are inserted throughout the condensate volume, and $j_{r}$ increases almost linearly with $r$. The scaffold flux at the surface of the core $r=r_{\textrm{c}}$ is set by the rate of core incorporation. If $q=0$, then $j_{r}=0$ at $r=r_{\textrm{c}}$ (figure 4; centre). In this case, scaffold molecules initially situated at the core surface remain at the core surface during condensate growth, and so cannot be stretched in the azimuthal direction ($B_{\theta\theta}|_{r=r_{\textrm{c}}}=1$). In the radial direction, however, the outwards flux of condensate material stretches the scaffold and induces large tensile strains. The scaffold flux resulting from the combination of both core and bulk incorporation is as expected; $j_{r}$ increases linearly far away from the core, but is non-zero at the core surface (figure 4; right). Correspondingly, we find large strains near to the core (which may be either tensile or compressive depending on the magnitude of $q$), and uniform tensile strains away from the core.
@@ II.7 Scaffold growth induces large strains near the core
II.8 Internal elasticity restricts condensate growth
To investigate the impact of internal elasticity on condensate growth, we now couple scaffold density to elastic deformation, allowing $\bar{\phi}_{S}$ to vary with stress as described by (8). Figure 5A shows that the influence of elastic stress on growth depends strongly on the rate of stress relaxation, quantified by $\tau$, compared to the rate of material incorporation. In the limit $k_{+}\ll\tau^{-1}$, strains relax sufficiently fast that condensate growth is unaffected by internal stresses. However, for $k_{+}\gg\tau^{-1}$ strain relaxation is slow compared to the rate of material incorporation and growth is significantly inhibited. Here, we keep $k_{+}$ and $q$ constant, implying that new material is incorporated into the condensate at the same rate regardless of the stress state. Consequently, elastic stresses have only a small effect on the rate of precursor material consumption (figure 5B). This result highlights that the restricted growth is driven by increased deformation of the scaffold, rather than by reaction kinetics. We explain the impact of elasticity on growth as follows: as new material is incorporated into the condensate, the scaffold volume fraction $\phi_{S}$, and hence the osmotic pressure, increase locally. To reduce osmotic pressure and obtain the preferred volume fraction $\phi_{S}^{0}$, the condensate expands, with solvent swelling the scaffold. This expansion stretches the scaffold, inducing a large tensile stress. This stress then acts against the osmotic pressure, resisting the increase in tension associated with further swelling. As a result, the condensate cannot expand fully, and remains at a smaller size, with $\bar{\phi}_{S}>\phi_{S}^{0}$. Over time, stress relaxation allows the osmotic pressure to drive further scaffold expansion, until eventually $\bar{\phi}_{S}=\phi_{S}^{0}$ and the condensate reaches its final (unstrained) volume. Faster stress relaxation (corresponding to a more liquid-like rheology) therefore facilitates faster condensate growth.
@@ II.8 Internal elasticity restricts condensate growth
II.9 Viscoelastic material properties impart mechanical strength
@@
III Centrosomes as growing viscoelastic materials
Centrosomes are membrane-less organelles that act as the main microtubule organising centre in metazoans. Centrosomes consist of a well-defined core (a pair of centrioles), which is surrounded by an amorphous, mesh-like structure known as the pericentriolar material (PCM). The PCM comprises a porous scaffold network permeated by both solvent and client molecules [21]. Recent experiments have shown that the PCM in C. elegans centrosomes exhibits a characteristic pore size [22], and displays a viscoelastic rheology [23]. During mitosis, centrosomes mature rapidly, with the PCM dramatically increasing in size. Incorporation of new scaffold material is driven by phosphorylation-dependent conversion of precursor molecules into an assembly-competent form within the pre-existing PCM [24, 25, 26]. As centrosomes grow, microtubules nucleate within the PCM, extending outwards into the surrounding cell. These microtubules utilise centrosomes as mechanical anchor points for transducing mechanical forces that organise the mitotic spindle [27]. Current models of centrosome structure, however, are unable to reconcile the twin requirements of rapid (fluid-like) PCM assembly and (solid-like) mechanical strength [9, 10]. To demonstrate the ability of our modelling framework to explain this dichotomy, we compare the predictions of our model to experimental data for centrosomes in C. elegans embryos, following the procedure detailed in Ref. [28].
@@ III Centrosomes as growing viscoelastic materials
Figure 6: (A) Phase space showing the quality of fit between simulation results and experimental growth curves for different elastic modulus $K$ and relaxation time $\tau$. The white curve separates parameter values for which an acceptable fit is possible, and for which it is not. (B) Normalised change in centrosome radius $\Delta R$ induced by a mitotic stress for various $K$ and $\tau$. The white curve separates parameter values for which scaffold expansion is acceptable, and for which it is not. (C) Combining results from panels (A) and (B) provides constraints on the material properties that allow both rapid growth and mechanical strength. (D) Comparison of model predictions (solid lines) and experimental data (scattered dots) for centrosome growth at various cell stages, assessed using GFP:SPD-5. For the experiments, $n=17$ embryos (one-cell stage), $18$ (two-cell), and $15$ (four-cell). Time is measured relative to cytokinesis onset of each division cycle. Values of material parameters indicated by green triangle in panel (A). Inset shows model predictions in the ‘unacceptable’ parameter regime (red square). (E) Depolymerisation of microtubules via nocodazole during both S-phase (purple) and at NEBD (green) inhibits scaffold growth compared to the control case (blue). Experiments were performed for control embryos ($n=14$ embryos), as well as embryos treated with nocodazole at S-phase ($n=9$) and at NEBD ($n=13$) of the first embryonic mitosis. Note slight reduction in PCM size upon extended treatment with nocodazole in line with model predictions. Time is measured relative to chromosome decondensation. (F) Proportion of total scaffold material that is incorporated at the centriole as a function of base incorporation rate $q_{0}$. Light to dark lines show simulations for increasing $\beta$ from $\beta=10^{-5}$ to $\beta=0.1$, which characterises how strongly reactions depend on strain. Other parameter values are the same as in figure 5. In panels (D) and (E), scattered dots show mean experimental values, and shaded areas show one standard deviation about the mean. We convert our results to dimensional units by setting $k_{-}=10^{-3}\,\textrm{s}$, $r_{\textrm{c}}=75\,\textrm{nm}$, and $\alpha=0.1\,\textrm{kPa}^{-1}$.
III.1 Liquid-like centrosomes permit rapid scaffold growth
@@
During centrosome maturation, the PCM scaffold undergoes rapid expansion as new scaffold proteins (e.g., SPD-5 in the case of C. elegans) are incorporated. However, scaffold incorporation also drives large strains within the PCM, which can restrict growth (figure 5). Comparing experimental growth curves for centrosomes from embryos at the 1, 2 and 4 cell stages with model predictions, we can constrain the mechanical parameters that allow sufficiently rapid growth (figure 6A). To determine permissible material properties, we fit our model to experimental data by varying $k_{+}$ and $\bar{\phi}$ for given values of the relaxation time $\tau$ and elastic modulus $K$. We keep the rate of centriolar incorporation $q$ constant throughout the fit, since it has a very limited effect on the overall growth curves. Quantifying the goodness of fit for each combination of material parameters by the mean squared error (figure 6A), we then identify ‘acceptable’ and ‘unacceptable’ regions of the parameter space for $\tau$ and $K$ spanning several orders of magnitude. We note that because elastic effects enter our model in combination with the phenomenological parameter $\alpha$, the dimensional values of $K$ presented here can be rescaled without affecting our general conclusions. For large values of $\tau$ and $K$ (solid-like rheology), we find that model predictions are incompatible with experimental data. In this parameter regime, a good fit between experiment and theory is not possible (figure 6D; inset), regardless of the fit parameters $k_{+}$ and $\bar{\phi}$, since elastic stresses limit the maximum rate of growth. On the other hand, for small values of $\tau$ and $K$ (liquid-like rheology), an excellent fit is possible (figure 6D; main).
III.2 Viscoelastic centrosomes can sustain mitotic forces
As well as exhibiting rapid growth during maturation, centrosomes must anchor large mechanical forces during mitosis. The action of mitotic forces is represented in our model by the external stress $\sigma^{*}$. Applying a mitotic force of $100\,\textrm{pN}$ [29, 30] to a fully grown centrosome, we then simulate the subsequent volumetric expansion of the scaffold over a time period of $100\,\textrm{s}$. For relatively liquid-like material properties (small $\tau$, $K$), we observe significant expansion of the scaffold, as quantified by the normalised change in radius $\Delta R$ (figure 6B). Such a large deformation may result in material failure and disassembly of the PCM. For more solid-like properties (large $\tau$, $K$), however, the scaffold is able to resist deformation and only expand by a small amount. As before, we now identify regions of the parameter space for which the simulated expansion is acceptable and for which it is not, using an expansion threshold of $1\%$. Combining the constraints identified in figures 6A-B, we can thus identify a parameter regime for which both rapid growth and mechanical strength are possible (figure 6C).
III.3 Microtubule-mediated stress drives scaffold expansion
To assess the impact of microtubule-mediated stresses on centrosome growth during maturation, we conducted experiments in which embryonic C. elegans is treated with nocodazole, which acts to depolymerise microtubules throughout the cell. This process can be simulated in our modelling framework by setting $\sigma^{*}=0$ at the point of nocodazole addition. Due to the finite time of nocodazole permeation, we enforce this change in $\sigma^{*}$ smoothly over a time period of $30\,\textrm{s}$.
Figure 6E shows centrosome volume as a function of time (measured relative to chromosome decondensation) for embryos treated with nocodazole both at nuclear envelope breakdown (NEBD) and during S-phase. For both treatment protocols, the final centrosome volume is significantly reduced in comparison to the control case. For centrosomes treated at NEBD, growth before treatment follows that of the control case. Once nocodazole is added, however, there is a sudden drop in centrosome volume. Notably, this decrease in volume does not coincide with a decrease in the total amount of scaffold protein in the PCM but instead results from an increase in scaffold density [28]. From a fit of model predictions with $\tau=100\,\textrm{s}$ to the experimental growth curves, we estimate the ratio of microtubule-mediated stress $\sigma^{*}$ to the elastic modulus $K$, which is on the order of $\sigma^{*}/K\sim 30$. For centrosomes treated during S-phase (i.e., before maturation has begun), centrosome volume is always smaller than for the control case. Additionally, the final centrosome volume is also smaller than for centrosomes treated during NEBD. We hypothesise that this result is due to NEBD-treated centrosomes initially growing under the expansion of the microtubule-mediated stress, thus leading to more bulk incorporation.
III.4 Stress-dependent reactions promote bulk incorporation